OpenStudy (anonymous):

A bullet is fired horizontally from the top of a building with a muzzle velocity of 150 m/s.A similar bullet dropped from the top of the same building, takes 4 sec to reach the ground. How far forward does the first bullet go before it hits the ground?

OpenStudy (anonymous):

since the bullet which is dropped takes 4 seconds u can calculate the height by using the formula s = ut + 1/2 at^2 her u = 0 s = 1/2 * 10 * 16 = 80 m similarly if the bullet is porjected horizontally also it will take 4 seconds to reach the ground. there is no acceleration in horizontal diretion . so distance travelled in horizontal direction = time * initial component = 4 * 150 =600 m/s

OpenStudy (anonymous):

hey can anyone tell me whether my answer is correct?

OpenStudy (anonymous):

Discounting wind resistance and the like, yes, your answer is correct...though it should read "600 meters," not "600 m/s." Basically, regardless of horizontal movement, gravity affects all objects the same. So, a dropped bullet hits the ground in the same amount of time as a bullet fired perpendicularly to the ground from the same height. So, if it takes four seconds for the dropped bullet to hit, it takes four seconds for the fired bullet to hit, too. If the fired bullet continues to travel 150 meters per second, in four seconds it will have traveled 600 horizontal feet.